OpenStudy (anonymous):

Determine if 3 * 4^1/6 is the simpliest form of 29161/6 . If not, tell why not and write the expression in simplest form.

OpenStudy (anonymous):

the question doesn't really make any sense the way it is written how could \[3\times 4^{\frac{1}{6}}\] be equal to \[\frac{29161}{6}\]?